# Reclaiming “Open AI" - AI Model Serving Can Be Open Access, Yet Monetizable and Loyal

### v3

@@ 2.1 The OML Protocol and Design Space
    Ownership Key ($k_{own}\in\mathcal{K}_{own}$): A secret key, unique to $\mathcal{O}$, essential for the permissioning scheme $\sigma$. It can be either cryptographic or AI-native, depending on the specific scheme.
@@
    Permission Scheme ($\sigma,h$): The functions $h:\mathcal{X}\to\mathcal{H}$ and $\sigma:\mathcal{H}\times\mathcal{K}_{own}\to\mathcal{P}$ that define how permission tokens are generated and bound to inputs. The security of this scheme (e.g., unforgeability of $p_{x}$) is paramount.
@@
    OMLization Process ($OMLize$): A transformation $M_{oml}\leftarrow OMLize(M,\text{params}_{\sigma,h})$. This process embeds the logic within $M_{oml}$ to verify $p_{x}$ and conditions its correct execution on this verification. $\text{params}_{\sigma,h}$ might include public parameters of the permission scheme (e.g., a public verification key $vk_{own}$ corresponding to $k_{own}$).
2.2 Core Properties of the OML Format
Let $M:\mathcal{X}\to\mathcal{Y}$ be a plain-text AI model that we want to protect, where $\mathcal{X}$ is the input space and $\mathcal{Y}$ is the output space. The OML cryptographic primitive transforms $M$ into an OML-formatted model, $M_{oml}$, designed to enable its distribution while protecting the intellectual property (IP) rights of the model owner, denoted $\mathcal{O}$. This balance ensures that AI models can be monetized effectively and remain loyal to $\mathcal{O}$.
Property 2.1 (Open-access (O)).
The Open-access (O) property signifies that $M_{oml}$ is openly accessible to everyone which provides the following specific operational benefits:
  • (a)
    Local Hosting and Execution: $M_{oml}$ is available for download (e.g., via a distribution channel $\mathcal{D}$), allowing users to host and execute it on their local infrastructure. This facilitates workflow optimization, can enhance data privacy for the user by keeping inputs local, and ensures consistent service quality independent of $\mathcal{O}$’s serving infrastructure.
  • (b)
    Local Enhancements: Users can, upon approval from the model owner $\mathcal{O}$, apply standard model improvement techniques locally to their copy of $M_{oml}$ (e.g., fine-tuning with local data, retrieval-augmented generation).
  • (c)
    Integrity and Post-Publication Immutability: The open accessibility of an $M_{oml}$ artifact provides strong assurances of its integrity. Specifically, once $M_{oml}$ is published and distributed, the model owner $\mathcal{O}$ cannot unilaterally alter that particular instance. This immutability by the owner ensures that users are interacting with the model as it was originally released, fostering trust in the specific version they possess.
Property 2.2 (Monetizable (M)).
The Monetizable (M) property ensures that an OML-formatted model $M_{oml}$ enables its owner $\mathcal{O}$ to derive economic value for each usage through a fine-grained, per-instance authorization protocol. The core mechanisms are detailed as follows:
  • (a)
    Granular Authorization: Access to the utility of $M_{oml}$ for a specific input $x\in\mathcal{X}$ requires a valid permission token $p_{x}$. This token is issued or verifiable by $\mathcal{O}$ and is essential for implementing flexible, pay-per-use billing models (e.g., per-input or per-token charges). This granular approach contrasts sharply with typical open-weight distributions where monetization strategies are often restricted to coarse-grained, model-level access (e.g., a one-time purchase for full model download or a general license for unlimited use).
  • (b)
    Conditional Value Proposition: Let $U(M,x)$ denote the utility derived from the output $M(x)$ of the original plain-text model. The OML model $M_{oml}$ ensures that a user $\mathcal{U}$ can only achieve utility comparable to $U(M,x)$ by adhering to $\mathcal{O}$’s terms for that specific usage instance. This adherence centrally involves acquiring and utilizing the valid permission token $p_{x}$, as unauthorized attempts to use $M_{oml}$ should not yield the desired utility.
Property 2.3 (Loyal (L)).
Loyalty ensures that an OML-formatted model $M_{oml}$ strictly adheres to its owner $\mathcal{O}$’s operational terms, primarily through the pre-hoc enforcement of authorization. This pre-hoc control is a crucial mechanism for addressing AI safety and alignment concerns by enabling owner-defined governance over the model’s use.
  • (a)
    Proof-of-Ownership: The owner $\mathcal{O}$ possesses a secret key $k_{own}\in\mathcal{K}_{own}$, which is fundamental to the generation or verification of permission tokens $p_{x}$ required for authorized use. The secret key can be either AI-native or cryptographic depending on the specific scheme.
  • (b)
    Pre-hoc Authorization Enforcement: $M_{oml}$ is engineered such that it will not produce a correct or high-utility output for an input $x$ unless a valid permission token $p_{x}$ (derived using $k_{own}$) is provided as part of its input. This authorization must occur before $M_{oml}$ completes its useful computation. Such enforcement is critical for high-value models and for scenarios requiring diligent control over AI safety and alignment, as it proactively prevents unauthorized or non-compliant content generation.
  • (c)
    Controlled Openness: While $M_{oml}$ itself might be openly accessible (as per Property 2.1), its correct and intended functioning is inextricably tied to $\mathcal{O}$’s authorization mechanisms. This owner-centric control can be designed to be auditable if the authorization protocol itself is transparent (e.g., through rules managed by a smart contract).
These three properties collectively define the desired behavior of an OML model. Crucially, the OML format must guarantee that unauthorized users cannot circumvent these IP rights and access controls to gain unpaid or unrestricted access to $M$’s capabilities through $M_{oml}$ without incurring prohibitive computational expense.
@@ 2.3 Security Guarantees and Adversarial Model
The primary security objective of the OML primitive is to enforce Properties 2.2 (Monetizability) and 2.3 (Loyalty), even when an adversary $\mathcal{A}$ has white-box access to $M_{oml}$.
Adversary Definition and Capabilities.
An adversary $\mathcal{A}$ is a computationally bounded algorithm whose goal is to circumvent the OML protections. $\mathcal{A}$ is assumed to possess:
  • Full white-box access to $M_{oml}$ (i.e., its architecture and parameters).
  • Access to an oracle $\mathcal{P}_{\mathcal{O}}(\cdot)$ that, for a chosen input $x_{i}\in\mathcal{X}$, returns the legitimate permission token $p_{x_{i}}=\sigma(h(x_{i}),k_{own})$. $\mathcal{A}$ can make up to $N$ queries to $\mathcal{P}_{\mathcal{O}}(\cdot)$, potentially at some cost per query. From these, $\mathcal{A}$ obtains a set of authorized input-token-output tuples $\mathcal{D}_{known}=\{(x_{i},p_{x_{i}},y_{i})\}_{i=1}^{N}$, where $y_{i}=M_{oml}(x_{i},p_{x_{i}})$.
Core Security Requirements (Hardness Properties).
We formally define the security of an OML scheme as follows. A secure OML scheme should satisfy both requirements 2.1 and 2.2.
@@
It must be computationally infeasible for an adversary $\mathcal{A}$, given $M_{oml}$ and $\mathcal{D}_{known}$, to construct a standalone model $M^{\prime}$ that effectively replicates the functionality of the original model $M$ without requiring the OML permission mechanism. Consider the following experiment, $Expt_{\mathcal{A}}^{ME}$:
@@
    $\mathcal{A}$ is given $M_{oml}$ and oracle access to $\mathcal{P}_{\mathcal{O}}(\cdot)$ for $N$ queries.
@@
    $\mathcal{A}$ outputs a model $M^{\prime}$.
@@
    $\mathcal{A}$ wins if, for a fresh input $x^{*}\sim\mathcal{D}_{\mathcal{X}}$ (drawn from the target distribution $\mathcal{D}_{\mathcal{X}}$, and $x^{*}\notin\{x_{i}\}_{i=1}^{N}$), $d(M^{\prime}(x^{*}),M(x^{*}))\leq\epsilon_{utility}$ with high probability.
The OML scheme is $(t,N,\epsilon_{ME})$-extraction resistant if for any adversary $\mathcal{A}$ running in time complexity at most $t$ and making $N$ queries, $Pr[Expt_{\mathcal{A}}^{ME}\text{ wins}]\leq\epsilon_{ME}$.
Requirement 2.2 (Permission Forgery Resistance).
It must be computationally infeasible for an adversary $\mathcal{A}$, given $M_{oml}$ and $\mathcal{D}_{known}$, to generate a valid permission token $p^{*}_{x^{*}}$ for a new input $x^{*}\notin\{x_{i}\}_{i=1}^{N}$ that allows $M_{oml}$ to produce a correct output for $x^{*}$. Consider the following experiment, $Expt_{\mathcal{A}}^{PF}$:
  1. 1.
    $\mathcal{A}$ is given $M_{oml}$ and oracle access to $\mathcal{P}_{\mathcal{O}}(\cdot)$ for $N$ queries.
  2. 2.
    A challenge input $x^{*}\sim\mathcal{D}_{\mathcal{X}}$ ($x^{*}\notin\{x_{i}\}_{i=1}^{N}$) is given to $\mathcal{A}$.
  3. 3.
    $\mathcal{A}$ outputs a permission token $p^{*}_{x^{*}}$.
  4. 4.
    $\mathcal{A}$ wins if $d(M_{oml}(x^{*},p^{*}_{x^{*}}),M(x^{*}))\leq\epsilon_{utility}$.
The OML scheme is $(t,N,\epsilon_{PF})$-forgery resistant if for any adversary $\mathcal{A}$ running in time at most $t$ and making $N$ queries, $Pr[Expt_{\mathcal{A}}^{PF}\text{ wins}]\leq\epsilon_{PF}$. This implies that $\mathcal{A}$ cannot recover $k_{own}$ or find a practical way to compute $\sigma(h(x^{*}),k_{own})$ without $k_{own}$.
Here, $\epsilon_{ME}$ and $\epsilon_{PF}$ should be negligible functions of the security parameter. Satisfying both requirements is essential for robustly ensuring loyalty and monetizability.
Threat Model: Potential Attack Strategies. Adversaries may employ several strategies to violate the above security requirements:
  1. 1.
    Verification Bypass or Removal: $\mathcal{A}$ analyzes $M_{oml}$ (e.g., using techniques from neural network surgery [85, 36] or program analysis) to identify and excise or neutralize the embedded permission verification logic. A successful attack would yield a model $M^{\prime}$ that behaves like $M$ but without requiring valid tokens, thus violating Requirement 2.1.
  2. 2.
    Verification Tampering: $\mathcal{A}$ attempts to manipulate the internal state or execution flow of $M_{oml}$ at runtime to force the verification step to succeed illicitly. This could lead to violating Requirement 2.2 for specific inputs if a generic bypass is not achieved.
  3. 3.
    Permission Forgery or Counterfeiting: $\mathcal{A}$ aims to produce valid $(x^{*},p^{*}_{x^{*}})$ pairs without legitimate access to $\mathcal{P}_{\mathcal{O}}(\cdot)$ for $x^{*}$. This includes:
    • (a)
      Directly crafting a $p^{*}_{x^{*}}$ that passes verification by exploiting weaknesses in $\sigma$ or its implementation within $M_{oml}$.
    • (b)
      Recovering $k_{own}$ or learning a functional equivalent of $\sigma(\cdot,k_{own})$ from $\mathcal{D}_{known}$.
    • (c)
      Training a surrogate model $M_{surrogate}(x)\approx p_{x}$ using $\mathcal{D}_{known}$ as training data.
Addressing Naive Constructions. A naive approach, such as $M_{oml}(x,p_{x}):=\text{If }Verify_{vk_{own}}(h(x),p_{x})\text{ then }M(x)\text{ else }\perp$, where $Verify$ is a standard digital signature verification algorithm and $p_{x}=Sign_{k_{own}}(h(x))$, is insecure. If $Verify$ and $M$ are loosely coupled components within $M_{oml}$, an adversary with white-box access can often easily identify and remove the $Verify$ call, effectively obtaining $M$. The $OMLize$ process must therefore deeply entangle the model’s computation with the verification logic.
The overarching goal of an OML primitive is to ensure that the computational cost for $\mathcal{A}$ to succeed in any of these attacks significantly outweighs the value derived, or that the success probability is negligibly small. This requires a sophisticated combination of cryptographic techniques, secure programming, and potentially obfuscation or machine learning-based defenses. By establishing these hardness properties, OML aims to create a fair and balanced ecosystem for AI models, protecting the rights of model owners without unduly restricting legitimate users.
3 Road to OML: Preliminary Solutions and Frameworks
This section outlines preliminary solutions and frameworks toward achieving robust Open-access, Monetizable, and Loyal (OML) model serving. We first explore theoretical OML constructions. We then discuss a practical instantiation, OML 1.0, based on AI-native model fingerprinting, alongside a proposed end-to-end framework for hosting and managing OML-formatted models. Detailed analyses for these topics are in Appendices A, B, and C.
3.1 OML in Theory - Canonical OML Construction Ideas
Various approaches can realize OML formatting, differing in their security guarantees and performance implications. These constructions, presented broadly by increasing security strength, include:
  1. 1.
    Obfuscation [Software security]: Leverage multi-layered software obfuscation techniques to protect models by making them hard to analyze and modify.
  2. 2.
    Trusted Execution Environments (TEEs) [Hardware security]: Leverage hardware-enforced secure enclaves to protect models during execution on remote untrusted machines.
  3. 3.
    Cryptography [Provable security]: Leverage primitives like Fully Homomorphic Encryption (FHE) for the strongest and provable security guarantees, often with performance trade-offs.
  4. $\star$
    Melange – mixed OML: Combine the aforementioned methods flexibly to tailor security and performance to specific needs, applying different techniques to different model parts.
These construction-based OML candidates often come with efficiency overheads (especially for TEEs and advanced Cryptography), and their application can be strategic. They can serve as robust backup mechanisms for dispute resolution or challenge verification in hybrid systems that primarily rely on more efficient API-based black-box access, with the OML-protected model acting as the ultimate source of truth when users are unsatisfied with the provided output and challenges arise.
Detailed discussions of these approaches, including their respective formatting and verification phases, characteristic trade-offs, and a comprehensive comparative analysis, can be found in Appendix A.
3.2 OML in Practice: OML 1.0 with AI-Native Fingerprinting and Ecosystem Framework
As a first practical step towards OML, OML 1.0 utilizes AI-native cryptography, specifically model fingerprinting. This technique repurposes data poisoning methods to embed unique, secret (key, response) pairs into models. A key challenge in this domain is the design of fingerprints that are robust against removal or detection, even when an adversary has white-box access to the model and its entire inference process. The workflow of OML 1.0, optimistic OML, is shown in Figure 3.

### v4

@@ 2.1 Properties and Design Space
    Authorization: Users submit $h(x)$ to owner $\Pi_{\mathcal{O}}$; if approved, they receive $p_{x}$ and query $(x,p_{x})$.
@@
    Fidelity: $d(M_{oml}(x,p_{x}),M(x))\leq\epsilon_{\text{utility}}$, ensuring preservance of the model’s core capabilities.
@@
    Protection: For invalid $p$, $d(M_{oml}(x,p),M(x))>\epsilon_{\text{robust}}$ with $\epsilon_{\text{robust}}>\epsilon_{\text{utility}}$.
  4. 4.
    Overhead: $T(M_{oml},(x,p_{x}))\leq(1+\epsilon_{\text{overhead}})\,T(M,x)$, preserving practical deployability.
These four promises collectively define what we term the quality profile $(\epsilon_{\text{utility}},\epsilon_{\text{robust}},\epsilon_{\text{overhead}})$, a quantitative characterization of an OMLization’s effectiveness. And the instantiation requires careful design of three interconnected components that form the technical foundation of any OML construction listed below. We also depict a high-level OMLization process in Algorithm 1.
  • an ownership key $k_{\text{own}}$ (either cryptographic or AI-native);
  • a binding/permission token issuance mechanism $(h,\sigma)$, optionally exposing $vk_{\text{own}}$;
  • verifier entanglement, which couples authorization to critical computations so that the high-utility pathway is reachable only under valid authorization.
Algorithm 1 OMLize: Transforming Models into Controlled Artifacts
1:Input: Original model $M$, binding function $h$, token scheme $\sigma$, public parameters
2:Output: Controlled artifact $M_{\text{oml}}$
3:Step 1: Embed verifier $\alpha:\mathcal{X}\!\times\!\mathcal{P}\!\to\!\{0,1\}$ that validates tokens against input commitments
4:Step 2: Entangle $\alpha$ within $M$’s critical paths to construct $F$ such that:
5:  (i) Valid authorization: $\alpha(x,p_{x})=1\Rightarrow F(x,p_{x})\approx M(x)$
6:  (ii) Invalid tokens: $\alpha(x,p)\neq 1\Rightarrow F(x,p)$ yields degraded/noisy output
7:Step 3: Optionally expose $vk_{\text{own}}$ for public verification capability
8:return $M_{\text{oml}}(x,p)=F(x,p)$
@@ 2.2 Security Guarantees and Adversarial Model
The security of OML must be analyzed under the assumption of white-box access where adversaries can inspect, modify, and experiment with $M_{oml}$ arbitrarily. This threat model reflects the reality where OMLized models may be controlled by potentially adversarial users.
Adversary Model. We model adversaries as probabilistic polynomial-time (PPT) algorithms $\mathcal{A}$ with
  • Complete white-box access to $M_{oml}$, including all parameters and computation graphs
  • Oracle access to an authorization service $\Pi_{\mathcal{O}}$ for up to $N$ queries
  • The resulting knowledge base $\mathcal{D}_{known}=\{(x_{i},p_{x_{i}},y_{i})\}_{i=1}^{N}$ where $y_{i}=M_{oml}(x_{i},p_{x_{i}})$
Security Goal. Against such adversaries, two fundamental hardness properties should hold:
@@
In experiment $\mathrm{Expt}^{\mathrm{ME}}_{\mathcal{A}}$:
(1) $\mathcal{A}$ receives $M_{oml}$ and oracle access to $\mathcal{P}_{\mathcal{O}}$ for $N$ queries; (2) $\mathcal{A}$ outputs a stand-alone model $M^{\prime}$;
(3) a fresh $x^{*}\sim\mathcal{D}_{\mathcal{X}}$ is drawn with $x^{*}\notin\{x_{i}\}$; (4) $\mathcal{A}$ wins if $d(M^{\prime}(x^{*}),M(x^{*}))\leq\epsilon_{utility}$.
The scheme is $(t,N,\epsilon_{ME})$-extraction-resistant if every PPT $\mathcal{A}$ running in time $t$ wins with probability at most $\epsilon_{ME}(t,N)$. Informally, any adversary cannot replicate a functionally equivalent model that bypasses authorization within reasonable cost.
Requirement 2.2 (Permission Forgery Resistance).
In experiment $\mathrm{Expt}^{\mathrm{PF}}_{\mathcal{A}}$:
(1) $\mathcal{A}$ receives $M_{oml}$ and oracle access to $\mathcal{P}_{\mathcal{O}}$ for $N$ queries;
(2) a fresh $x^{*}\sim\mathcal{D}_{\mathcal{X}}$ is revealed with $x^{*}\notin\{x_{i}\}$;
(3) $\mathcal{A}$ outputs $p^{*}$; (4) $\mathcal{A}$ wins if $d(M_{oml}(x^{*},p^{*}),M(x^{*}))\leq\epsilon_{utility}$.
The scheme is $(t,N,\epsilon_{PF})$-forgery-resistant if every PPT $\mathcal{A}$ running in time $t$ wins with probability at most $\epsilon_{PF}(t,N)$. Informally, adversaries cannot generate valid tokens for unauthorized inputs.
These requirements must withstand a diverse threat landscape:
@@
    Surgical Extraction: Adversaries employ network surgery techniques [80, 32] to identify and excise authorization logic while preserving model functionality.
@@
    Runtime Manipulation: Fault injection or state tampering forces the internal verifier to accept invalid tokens, bypassing authorization checks without modifying the model itself.
@@
    Cryptanalytic Forgery: Adversaries attempt to forge valid tokens through: (a) exploiting implementation vulnerabilities in $\sigma$, (b) recovering the secret key $k_{own}$ from side channels, or (c) training surrogate functions $\hat{\sigma}:x\mapsto p_{x}$ using $\mathcal{D}_{known}$.
The Failure of Naive Approaches.
To illustrate why sophisticated entanglement is necessary, consider a naive wrapper design with a cryptographic digital signature scheme:
$M_{oml}(x,p)\ :=\ \begin{cases}M(x)&\text{if }\mathrm{Verify}_{vk_{own}}(h(x),p)=\text{true}\\ \perp&\text{otherwise}\end{cases}$
With white-box access, an attacker can trivially locate the conditional branch, remove the verification check, and recover the original model $M$. This vulnerability motivates our requirement for deep computational entanglement, i.e. the verifier must be so thoroughly integrated that removing it is tantamount to destroying the model’s learned representations.
2.3 Theoretical Foundations and Limits
In this subsection, we state three results that define the feasible region for OML: an upper bound that prevents over-claiming, a sufficient condition that anchors OML in standard hardness, and an operational constraint that links security to issuance policy. Proofs are deferred to App. A.
First, if an adversary controls the artifact and can issue unbounded authorized queries, information alone suffices to reconstruct the task mapping, and perfect protection is therefore unattainable.
Theorem 2.2 (Information-theoretic impossibility).
No OML scheme achieves perfect security against unbounded adversaries with unlimited oracle access.
Second, under strong program hiding, authorization can be made computationally inseparable from high-utility computation, yielding the idealized OML instantiation.
Theorem 2.3 (OML from indistinguishability obfuscation).
If indistinguishability obfuscation (iO) exists for the model class, then there is an OML construction satisfying extraction and forgery resistance (assuming unforgeability of $\sigma$).
Third, authorized answers facilitate extraction. Learning theory converts model complexity and accuracy tolerance into a concrete cap on such answers.
Theorem 2.4 (Query–security trade-off).
Let $\mathcal{H}\subseteq[0,1]^{\mathcal{X}}$ have pseudo-dimension $d$ and assume $M\in\mathcal{H}$ (realizable). If an adversary receives $N$ i.i.d. authorized pairs and returns an ERM under squared loss, then there exist constants $C,c>0$ such that
$N\ \geq\ C\,\frac{d+\log(1/\delta)}{\varepsilon^{2}}\ \Rightarrow\ \Pr\!\big[\mathbb{E}\!(\hat{h}(x)-M(x))^{2}\leq\varepsilon\big]\geq 1-\delta,$
and any OML deployment targeting $(\varepsilon,\delta)$ extraction resistance must enforce $N<c\,\frac{d+\log(1/\delta)}{\varepsilon^{2}}$.
Implications. Taken together, Theorems 2.22.4 delineate the design space that motivates our concrete methods in the next sections: (i) Absolute guarantees are unattainable, so OML must rely on computational hardness and economics; (ii) Verifier entanglement with cryptographic binding is the appropriate abstraction for practical surrogates of iO; and (iii) Policies by model owners (token issuance, batching, collateral) must enforce query budgets consistent with the learned trade-off above. The constructions that follow instantiate these principles with varying efficiency–security profiles.
3 Road to OML: From Principles to Deployable Mechanisms
The transformation from theoretical primitive to practical system requires navigating fundamental trade-offs between security guarantees, computational efficiency, and deployment constraints. This section bridges the formal OML framework developed in Sections 2.12.3 with concrete implementations. We present canonical constructions that embed cryptographic authorization check $\alpha(x,p)$ through different mechanisms for perfect fidelity and protection guarantees, analyze the security, assumptions, and tradeoffs among different methods, and then introduce OML 1.0, an immediate deployment pathway using AI-native fingerprinting with optimal additional overhead.
3.1 Canonical Constructions: Security-Performance Spectrum
In this subsection, we choose cryptographic $\alpha(x,p)$ for authorization check, and cryptography schemes guarantee that $\epsilon_{robust}=\text{maximal}$ and $\epsilon_{utility}=0$, both achieving the optimal configuration. The challenge of OML implementation lies in making the model’s high-utility computation path accessible exclusively when $\alpha(x,p)=1$ within a small overhead $\epsilon_{overhead}$, while ensuring this entanglement cannot be surgically removed even under white-box access, i.e. Security requirements 2.1 and 2.2 are satisfied. We present four archetypal approaches for embedding authorization logic, each offering distinct trade-offs between security guarantees and deployment costs.
Obfuscation (Software Security).
In this scheme, $\alpha$ is embedded through graph transformations and code hardening. Authorization checks integrate with residual connections, attention mixing, and normalization paths. Graph permutation, control flow flattening, and constant blinding increase removal difficulty. It offers near-optimal $\epsilon_{\text{overhead}}\approx 0$ as hardly any additional computation is introduced. This method is immediately deployable but doesn’t have any provable security guarantee, and is vulnerable to dedicated hackers who try to reverse-engineer the obfuscation process.
[…]
1:Input: model $M$, binding $h$, token scheme $\sigma$, compiler/obf params
2:Verifier injection: Synthesize $\alpha(x,p)$; weave gates into critical paths (e.g., attention/key/value mixing, residual scalars).
3:Utility shaping: Construct $F$ so that $\alpha(x,p_{x}){=}1\Rightarrow F(x,p_{x})\approx M(x)$; else $F$ diverts to low-utility basins (e.g., masked subspaces, biased heads).
4:Hardening: Apply graph randomization (permute blocks), control-flow flattening, dead-code sprinkling, and constant blinding on verifier features.
5:Build: Compile with aggressive inlining; invoke multi-pass obfuscation/toolchain hardening.
6:Publish: $M_{\text{oml}}(x,p)\!=\!F(x,p)$, optional $vk_{\text{own}}$.
TEE-Gated Execution (Hardware Security).
TEE (Trusted Execution Environments) isolate critical subgraph in attested enclave where $\alpha$ verification gates computation. Unauthorized queries terminate before reaching utility paths. It enjoys moderate $\epsilon_{\text{overhead}}$ as the only overhead comes from enclave transitions. TEEs are already production-ready on CPUs, and thus suitable for models with narrow critical cores, but large models with billions of parameters cannot be OMLized with TEEs as commercialized GPUs are not available yet. Also, this method assumes correct vendor implementation and side-channel mitigations, introducing extra layer of trust and vulnerability.
[…]
1:Input: model $M$, binding $h$, token scheme $\sigma$, enclave config
2:Packaging: Encrypt $M$ and verifier code with enclave-sealed keys; Provision $vk_{\text{own}}$ as a public parameter.
3:Attestation: Publish measurement of enclave binary; expose remote attestation endpoint to $\Pi_{\mathcal{O}}$.
4:Authorization path: Inside TEE, verify $\alpha(x,p)\!=\!1$ against $h(x)$ and $vk_{\text{own}}$; otherwise exit with noise/denial.
5:Execution: Only upon successful verification, decrypt weights on-device with the enclave-sealed secret key, run $M$; Always re-encrypt with the public key before exiting the enclave.
6:Publish: $M_{\text{oml}}$ as an attested service binary + policy manifest.
Cryptographic Encryption (Provable Security).
Fully homomorphic encryption (FHE) offers a clean construction for OML: inputs are encrypted under a public key, the model is compiled to an arithmetic circuit and evaluated homomorphically on ciphertexts, and only the owner, who holds the secret key, can decrypt the result. Authorization is then enforced by decryption control: the owner decrypts outputs only for inputs carrying valid permissions. Under standard hardness assumptions (e.g., LWE), FHE provides provable security without external assumptions, becoming the perfectly secure OML construction. However, FHE evaluation incurs large multiplicative overhead that scales with circuit depth and bootstrapping frequency (often $10^{3}\!-\!10^{5}\times$ on today’s workloads), while exact FHE schemes over rings (BGV/BFV) operate on integers and thus require quantization which may downgrade performance, making this approach infeasible for full-scale LLM inference today.
[…]
1:Input: base model $M$, input-binding $h$, token scheme $\sigma$, FHE parameters (scheme, depth, scale), quantization policy
2:Key generation (owner): $(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{FHE.KeyGen}(\text{params})$. Publish $\mathsf{pk}$; keep $\mathsf{sk}$ secret.
3:Model-to-circuit: Compile $M$ to an arithmetic circuit $C_{M}$ respecting FHE depth (e.g., polynomial activations, folded norms). Apply quantization if using exact integer FHE.
4:Parameter protection: Encrypt model weights: $\widetilde{W}\!\leftarrow\!\mathsf{FHE.Enc}(\mathsf{pk},W)$.
5:Authorization channel: Specify decryption policy: owner will decrypt outputs iff presented with a valid token $p_{x}=\sigma(h(x),k_{\text{own}})$ (and optional usage proof/commitment).
6:Publish artifact: $(C_{M},\widetilde{W},\mathsf{pk},vk_{\text{own}})$ as the OML service interface.
Melange Hybrid (adaptive composition).
The mechanisms above can be composed by component criticality: e.g., Protect a minimal control core (e.g., routing heads or safety gates) with a TEE or a compact cryptographic subgraph, and harden the surrounding layers with software obfuscation. This Melange design lets owners tune the quality profile: the runtime cost scales with the size of the isolated core ($\epsilon_{\text{overhead}}$ controllable). Assumptions are localized to each layer: hardware trust for the enclave, cryptographic hardness for the small protected circuit, and program-analysis resistance for the periphery, yielding a practical, adaptive path to higher assurance without forfeiting openness.
3.2 OML 1.0: AI–Native Fingerprinting for Accountable Open Distribution
We present OML 1.0, an efficient instantiation that achieves monetizability and accountable loyalty through AI–native fingerprints: secret (key, response) pairs embedded in $M$ so that authorized service can be verified ex post with high confidence. Unlike wrapper checks, fingerprints are learned behaviors distributed across representational pathways and survive typical serving conditions. OML 1.0 offers post-hoc (“next-day security”) with near–zero additional inference overhead.
Core mechanism. Let $\mathcal{K}_{\mathrm{fp}}=\{(k_{i},r_{i})\}_{i=1}^{n}$ be a secret set of fingerprints. During OMLization, we fine-tune $M$ so that querying with $k_{i}$ elicits response $r_{i}$ and preserves task utility on the deployment distribution $\mathcal{D}_{\mathcal{X}}$. Model hosts sign licenses that require them to report every usage to the platform which issues per-input permissions and logs authorized usage. Independent provers periodically query public endpoints with hidden $k_{i}$; a correct $r_{i}$ without a matching authorization record constitutes a verifiable license violation (collateral slashing), enforcing the signed license, as depicted in Figure 2.